OpenStudy (mnr134):

What is the slope of the line? A.-1/7 B.1/7 C.–1 D.1

OpenStudy (stormer15):

line?

OpenStudy (stormer15):

\(\dfrac{-3-4}{2-(-5)}=\dfrac{-3-4}{2+5}=?\)

OpenStudy (stormer15):

?

OpenStudy (mnr134):

so its b

OpenStudy (stormer15):

Did you solve what I put above?

OpenStudy (mnr134):

yes

OpenStudy (stormer15):

What did you get?

OpenStudy (mnr134):

1/7

OpenStudy (stormer15):

-3 - 4 = -7 2 + 5 = 7 -7/7 = ?

OpenStudy (mnr134):

so its either a or c